# Bidirectional Uncertainty-Aware Region Learning for Semi-Supervised Medical Image Segmentation

### v1

@@ 3.4 Certainty-aware Region Learning
Specifically, for an unlabeled data ${X}^{l}_{i}$, we use the teacher model ${F}_{s}(\theta_{s})$ trained from the labeled data to generate initial pseudo-labels $\mathbf{Y}^{u}$. Similar to the URL, we perform the same operation on the unlabeled data to obtain the uncertainty region and certainty region predicted by the model, and at the same time, generate the uncertainty mask $\mathbf{M}^{{}^{\prime}}$ according to the entropy threshold set in advance, and classify the reliable pseudo-labels into reliable pseudo-labels and unreliable pseudo-labels by the uncertainty mask $\mathbf{M}^{{}^{\prime}}$, we would like to reduce the learning intensity of the uncertainty regions in the CRL. Unlike the URL, since unlabeled data does not have accurate labels and relies only on model-generated pseudo-labels, a high threshold needs to be set to localize these high-uncertainty regions, and if the threshold is set too low, it may contain many high-uncertainty regions that can mislead the model. The final pseudo-labeling is processed by the following formula:
$\mathbf{Y}^{u}_{c}=\mathbf{Y}^{l}\odot\left(\mathbf{1}-\mathbf{M^{{}^{\prime}}}\right),$ (6)
$\mathbf{Y}^{u}_{uc}=\mathbf{Y}^{l}\odot\mathbf{M^{{}^{\prime}}},$
@@ 3.5 Loss Functions
$\mathcal{L}^{l}={L}_{seg}$ $\left(\mathbf{P}^{l},\mathbf{Y}^{l}_{uc}\right)\odot\mathbf{M}+$ (7)
$\alpha{L}_{seg}\left(\mathbf{P}^{l},\mathbf{Y}^{l}_{c}\right)\odot\left(\mathbf{1}-\mathbf{M}\right),$
$\mathcal{L}^{u}={L}_{seg}$ $\left(\mathbf{P}^{u},\mathbf{Y}^{u}_{c}\right)\odot\left(\mathbf{1}-\mathbf{M^{{}^{\prime}}}\right)+$ (8)
$\alpha{L}_{seg}\left(\mathbf{P}^{u},\mathbf{Y}^{u}_{uc}\right)\odot\mathbf{M^{{}^{\prime}}},$
@@
In each iteration, we update the parameters in the student network by stochastic gradient descent of the loss function, and the parameters of the teacher network by exponential moving average ($EMA$[18]:
@@
where $t$ and ${t-1}$ denote the current training step and its previous one, respectively, $\theta_{s}$ and $\theta_{t}$ are the weights of the student and teacher models at training step t. where $\lambda$ is the smoothing coefficient parameter, with a recommended value of $\min\left(t/(t+1),0.99\right)$, which is used to control the updating rate of the teacher weight.
Method Scans used Metrics
[…]
U-Net 3(5%) 0 47.83 37.01 31.16 12.62
U-Net 7(10%) 0 79.41 68.11 9.35 2.70
U-Net 70(All) 0 91.44 84.59 4.30 0.99
MT 3(5%) 67(95%) 51.20 40.06 21.27 6.92
SASSNet 57.77 46.14 20.05 6.06
DTC 56.90 45.67 23.36 7.39
URPC 55.87 44.64 13.60 3.74
MC-Net 62.85 52.29 7.62 2.33
SS-Net 65.83 55.38 6.67 2.28
BCP 87.59 78.67 1.90 0.67
Ours(MT) 63.64 51.46 20.32 6.57
Ours(BCP) 88.42 79.83 1.82 0.64
MT 7(10%) 63(90%) 79.65 68.06 12.87 3.99
SASSNet 84.50 74.34 5.42 1.86
DTC 84.29 73.92 12.81 4.01
URPC 83.10 72.41 4.84 1.53
MC-Net 86.44 77.04 5.50 1.84
SS-Net 86.78 77.67 6.07 1.40
BCP 88.84 80.62 3.98 1.17
Ours(MT) 82.19 71.36 10.04 2.66
Ours(BCP) 90.40 83.01 1.63 0.41
Method Scans used Metrics
[…]
V-Net 4(5%) 0 52.55 39.60 47.05 9.87
V-Net 8(10%) 0 82.74 71.72 13.35 3.26
V-Net 80(All) 0 91.47 84.36 5.48 1.51
MT 4(5%) 76(95%) 81.74 69.93 16.30 4.99
SASSNet 81.60 69.63 16.16 3.58
DTC 81.25 69.33 14.90 3.99
URPC 82.48 71.35 14.65 3.65
MC-Net 83.59 72.36 14.07 2.70
SS-Net 86.33 76.15 9.97 2.31
BCP 88.02 78.72 7.90 2.15
Ours(MT) 82.88 71.40 11.11 3.18
Ours(BCP) 89.10 80.46 7.72 1.98
MT 8(10%) 72(90%) 86.81 76.93 10.07 2.49
SASSNet 87.54 78.05 9.84 2.59
DTC 87.51 78.17 8.23 2.36
URPC 86.92 77.03 11.13 2.28
MC-Net 87.62 78.25 10.03 1.82
SS-Net 88.55 79.62 7.49 1.90
BCP 89.62 81.31 6.81 1.76
Ours(MT) 87.95 78.77 8.54 2.11
Ours(BCP) 90.21 82.26 6.27 1.63
@@ 4.2 Evaluation Metrics
We choose four evaluation metrics: Dice Score (%), Jaccard Score (%), 95% Hausdorff Distance (95HD) in voxel and Average Surface Distance (ASD) in voxel. Dice [35] and Jaccard measure overlap percentages between two object regions, while ASD calculates the average boundary distance, and 95HD computes the closest point distance between them.
@@ 4.3 Implementation Details
In our experiments, we set $\alpha=0.5$. We performed all experiments on a device with an NVIDIA 3090 GPU with a fixed random seed. Due to the different feature representations and data structures of 2D and 3D medical image datasets, we conducted experiments on the ACDC dataset and the LA dataset. Specifically, in the ACDC dataset, we set the threshold $\mu$ to the 95th percentile of the entropy matrix predicted by the model each time during the supervision phase and the threshold $\mu$ to the 99th percentile during the self-training phase. In the LA dataset, the threshold $\mu$ was set to the 95th percentile of the entropy matrix in both cases. All other settings follow the default settings with the original Mean-Teacher and BCP.
@@ 4.4 Comparison with Sate-of-the-Art Methods
ACDC dataset
Table 1 shows the average performance of the four types of segmentation results on the ACDC dataset, with labeling ratios of 5% and 10%, respectively. After incorporating our methods into MT and BCP (i.e., Ours(MT) and Ours(BCP)), they showed significant improvements, which highlights the flexibility and scalability of our methods. Specifically, with a 5% labeling ratio, Ours(MT) achieves a Dice score that is 12.4% higher than the original MT. Ours(BCP) shows a Dice score improvement of 0.8% over the original BCP. Similarly, with a 10% labeling ratio, Ours(MT) achieves a Dice score that is 2.5% higher than the original MT, while Ours(BCP) achieves a Dice score improvement of 1.56% over the original BCP. These improvements in Dice scores, along with corresponding gains in Jaccard indices, 95HD, and ASD, demonstrate the effectiveness of our semi-supervised learning strategy in enhancing segmentation accuracy with limited labeled data.
Fig. 3 gives some qualitative results where our method effectively suppresses the regions that are inaccurately segmented by the Mean Teacher and BCP methods.
LA dataset
Following SS-Net, we conducted experiments with 5% and 10% labeled data. We compared our approach with MT, SASSNet [26], DTC [27], URPC [36], MC-Net [37], SS-Net [38] and BCP[39]. As shown in Table 2, our approach surpasses all other methods. Specifically, compared to the baseline, our method shows some improvement in model performance.
Fig. 4 gives some qualitative results where our method is effective in segmenting difficult boundary regions.
@@
Figure 3: Visualization of segmentation results on ACDC dataset with 10% labeled data. (a) Ground-truth. (b) Mean Teacher results. (c) Ours (Mean Teacher ) results. (d) BCP results. (e) Ours (BCP) results. Best viewed in color on the screen.
@@
Figure 4: Visualization of segmentation results on LA dataset with 10% labeled data. (a) Ground-truth. (b) Mean Teacher results. (c) Ours (Mean Teacher ) results. (d) BCP results. (e) Ours (BCP) results. Best viewed in color on the screen.
@@ 4.5 Ablation Studies
We conducted an ablation study to show the impact of each component on our network. This included the effectiveness of the bidirectional region learning strategy, the regional learning approach, uncertainty threshold $\mu$, and the impact of removing unreliable prediction regions on model learning.
4.5.1 Effectiveness of Each Module in Bidirectional Region Learning
As shown in Table 3, to validate the bidirectional region learning strategies that we employ in both phases, we explore the following approaches under Ours(BCP) 10% labeled data. We demonstrate the effectiveness of each module in dual-phase region learning by progressively adding URL and CRL, when URL was added, the Dice score increased to 89.45%, indicating that focusing on high-uncertainty regions in labeled data helps the model better learn from complex areas and transfer the learned knowledge to the training of unlabeled data. When CRL was added, the Dice score rose to 90.20%, showing that focusing on low-uncertainty regions in unlabeled data and reducing the learning weights helps mitigate the impact of erroneous pseudo-labels. By incorporating both URL and CRL into the baseline, the model has the best result, reaching 90.40% Dice and surpassing the baseline in 1.56%. These results indicate that the two modules are complementary during the training process, the model not only gains a deeper understanding of complex regions but also maintains stability and accuracy in the training of unlabeled data, thereby improving the overall performance of the model.
Table 3: Effectiveness of URL and CRL modules. “base” means baseline is the BCP.
Base URL CRL Metrics
Dice Jaccard 95HD ASD
@@ 4.5.1 Effectiveness of Each Module in Bidirectional Region Learning
Table 4: Impact of regional learning approach on model learning.
Region Metrics
Labeled Unlabeled Dice Jaccard 95HD ASD
@@ 4.5.2 Regional Learning Strategies
To evaluate the effectiveness of the URL and CRL modules on labeled and unlabeled data, under 10% labeled data based on Ours (BCP), we conducted different combination experiments with both modules on labeled and unlabeled data, as shown in Table 4. Through these combination strategies, we were able to analyze the effects of applying the two modules on different types of data. The experimental results indicate that all combination strategies brought significant performance improvements, demonstrating the effectiveness of the URL and CRL modules. The optimal result was achieved by applying URL on labeled data and CRL on unlabeled data. This result confirms our hypothesis: for labeled data, we should focus on learning uncertain regions, as accurately labeled data can provide clear guidance, helping the model better understand complex features and enhance its predictive ability. Conversely, for unlabeled data, where label guidance cannot be directly utilized, focusing on learning the deterministic regions of the model’s predictions can more effectively reduce the interference of pseudo-label noise, thereby improving the model’s robustness and generalization capability.
@@
Table 5: Model impact of uncertainty threshold $\mu$ settings in URL and CRL.
URL CRL Metrics
Dice Jaccard 95HD ASD
@@
Method Metrics
Dice Jaccard 95HD ASD

### v2

@@ 3.4 Certainty-aware Region Learning
Specifically, for an unlabeled data $\mathbf{X}^{l}_{i}$, we use the teacher model ${F}_{s}(\theta_{s})$ trained from the labeled data to generate initial pseudo-labels $\mathbf{Y}^{u}$. Similar to the URL, we perform the same operation on the unlabeled data to obtain the uncertainty region and certainty region predicted by the model, and at the same time, generate the uncertainty mask $\mathbf{M}^{\prime}$ according to the entropy threshold set in advance, and classify the pseudo-labels into reliable pseudo-labels and unreliable pseudo-labels by the uncertainty mask $\mathbf{M}^{\prime}$, we would like to reduce the learning intensity of the uncertainty regions in the CRL. It should be noted that after the training of the URL, the prediction results of the model are gradually stabilized, and the uncertain regions show a tendency of concentration and convergence, so we set a higher threshold in the CRL to accurately locate the uncertain regions. If the threshold is too low, some regions accurately predicted by the model may be incorrectly labeled as uncertain regions, thus affecting the training effect. The following formula processes the final pseudo-labeling:
$\mathbf{Y}^{u}_{c}=\mathbf{Y}^{u}\odot\left(\mathbf{1}-\mathbf{M^{\prime}}\right),$ (6)
$\mathbf{Y}^{u}_{uc}=\mathbf{Y}^{u}\odot\mathbf{M^{\prime}},$
@@ 3.5 Loss Functions
$\mathcal{L}^{l}$ $={L}_{seg}\left(\mathbf{P}^{l},\mathbf{Y}^{l}_{uc}\right)\odot\mathbf{M}+\alpha{L}_{seg}\left(\mathbf{P}^{l},\mathbf{Y}^{l}_{c}\right)\odot\left(\mathbf{1}-\mathbf{M}\right),$ (7)
$\mathcal{L}^{u}$ $={L}_{seg}\left(\mathbf{P}^{u},\mathbf{Y}^{u}_{c}\right)\odot\left(\mathbf{1}-\mathbf{M^{\prime}}\right)+\alpha{L}_{seg}\left(\mathbf{P}^{u},\mathbf{Y}^{u}_{uc}\right)\odot\mathbf{M^{\prime}},$ (8)
@@
In each iteration, we update the parameters in the student network by stochastic gradient descent of the loss function, and the parameters of the teacher network by exponential moving average (EMAEMA[19]:
@@
where $t$ and t1{t-1} denote the current training step and its previous one, respectively, $\theta_{s}$ and $\theta_{t}$ are the weights of the student and teacher models at training step t. where $\lambda$ is the smoothing coefficient parameter, with a recommended value of $\min\left(t/(t+1),0.99\right)$, which is used to control the updating rate of the teacher weight.
@@ 4.1.2 LA Dataset
4.1.3 PROMISE12 Dataset
The PROMISE12 dataset [43] was made available for the MICCAI 2012 prostate segmentation challenge. Magnetic Resonance (MR) images (T2-weighted) of 50 patients with various diseases were acquired at different locations with several MRI vendors and scanning protocols. All 3D scans are converted into 2D slices for segmentation tasks.
@@ 4.2 Evaluation Metrics
We choose four evaluation metrics: Dice Score (%), Jaccard Score (%), 95% Hausdorff Distance (95HD) in voxel and Average Surface Distance (ASD) in voxel. Dice [44] and Jaccard measure overlap percentages between two object regions, while ASD calculates the average boundary distance, and 95HD computes the closest point distance between them. The following are the formulas for these evaluation metrics:
[…]
$\text{95HD}=\max\{\sup_{a\in A}\inf_{b\in B}|a-b|,\sup_{b\in B}\inf_{a\in A}|b-a|\},$
$\text{ASD}=\frac{\sum_{a\in A}\inf_{b\in B}|a-b|+\sum_{b\in B}\inf_{a\in A}|b-a|}{|A|+|B|}.$
In these formulas, $A$ and $B$ represent the predicted and true label regions of the image, respectively. The terms $\sup_{a\in A}\inf_{b\in B}|a-b|$ and $\sup_{b\in B}\inf_{a\in A}|b-a|$ denote the maximum of the minimum distances from any point in set $A$ to the closest point in set $B$ and vice versa. By using these evaluation metrics, we can comprehensively assess the model’s performance across different dimensions, including overlap regions and boundary distances, allowing for a more accurate evaluation of the model’s effectiveness.
[…]
U-Net 5% 0% 47.83 37.01 31.16 12.62
U-Net 10% 0% 79.41 68.11 9.35 2.70
U-Net 100% 0% 91.44 84.59 4.30 0.99
MT 5% 95% 51.20 40.06 21.27 6.92
SASSNet 57.77 46.14 20.05 6.06
DTC 56.90 45.67 23.36 7.39
URPC 55.87 44.64 13.60 3.74
MC-Net 62.85 52.29 7.62 2.33
$S$-Net 65.83 55.38 6.67 2.28
BCP 87.59 78.67 1.90 0.67
GA-BAP 88.24 79.60 3.91 1.11
ABD 88.96 80.70 1.57 0.52
Ours(MT) 63.64 51.46 20.32 6.57
Ours(BCP) 88.42 79.83 1.82 0.64
Ours(ABD) 89.73 81.92 1.86 0.58
MT 10% 90% 79.65 68.06 12.87 3.99
SASSNet 84.50 74.34 5.42 1.86
DTC 84.29 73.92 12.81 4.01
URPC 83.10 72.41 4.84 1.53
MC-Net 86.44 77.04 5.50 1.84
$S$-Net 86.78 77.67 6.07 1.40
BCP 88.84 80.62 3.98 1.17
GA-BCP 89.31 81.27 3.32 1.01
ABD 89.81 81.95 1.46 0.49
Ours(MT) 82.19 71.36 10.04 2.66
Ours(BCP) 90.40 83.01 1.63 0.41
Ours(ABD) 90.18 82.72 2.27 0.69
@@ 4.3 Implementation Details
We selected several baseline models for our experiments: Mean Teacher (MT) [42], BCP [32], and ABD [33]. MT is a basic teacher-student network. BCP builds upon this by introducing a bidirectional copy-paste strategy for labeled and unlabeled data to generate mixed input data. ABD designs an adaptive bidirectional displacement strategy based on prediction confidence to generate new samples for retraining. To verify the effectiveness of our training strategy, we apply our method to these approaches to explore its adaptability to different baseline models and assess the performance improvement it brings. For ABD, we used only one teacher-student network for training in our experiments.
In the experiment, we set the default value of the learning weight to $\alpha=0.5$. We performed all experiments on an NVIDIA 3090 GPU device with a fixed random seed. Due to the differences between 2D and 3D medical image datasets, our experimental settings are also adjusted accordingly. Specifically, for the 2D ACDC and PROMISE12 datasets, the input size is cropped to $256\times 256$, we set the threshold $\mu$ to the 95th percentile of the entropy matrix predicted by the model each time during the supervision phase and the threshold $\mu$ to the 99th percentile during the self-training phase. For the 3D LA dataset, the input size is cropped to $112\times 112\times 80$, the threshold $\mu$ was set to the 95th percentile of the entropy matrix in both cases. All other settings follow the default settings with the original method.
[…]
V-Net 5% 0% 52.55 39.60 47.05 9.87
V-Net 10% 0% 82.74 71.72 13.35 3.26
V-Net 100% 0% 91.47 84.36 5.48 1.51
MT 5% 95% 81.74 69.93 16.30 4.99
SASSNet 81.60 69.63 16.16 3.58
DTC 81.25 69.33 14.90 3.99
URPC 82.48 71.35 14.65 3.65
MC-Net 83.59 72.36 14.07 2.70
$S$-Net 86.33 76.15 9.97 2.31
BCP 88.02 78.72 7.90 2.15
AllSpark 87.99 78.83 7.44 2.10
Ours(MT) 82.88 71.40 11.11 3.18
Ours(BCP) 89.10 80.46 7.72 1.98
MT 10% 90% 86.81 76.93 10.07 2.49
SASSNet 87.54 78.05 9.84 2.59
DTC 87.51 78.17 8.23 2.36
URPC 86.92 77.03 11.13 2.28
MC-Net 87.62 78.25 10.03 1.82
$S$-Net 88.55 79.62 7.49 1.90
BCP 89.62 81.31 6.81 1.76
AllSpark 88.74 80.54 7.06 1.82
Ours(MT) 87.95 78.77 8.54 2.11
Ours(BCP) 90.21 82.26 6.27 1.63
@@ 4.4 Comparison with Sate-of-the-Art Methods
4.4.1 ACDC dataset
To assess the generalization capability of our model across different domains, we conducted additional comparative experiments on the ACDC dataset with labeling ratios of 5% and 10%.. We present a comparative analysis of our framework against various competing methods: MT [42], SASSNet [29], DTC [30], URPC [41], MC-Net [45], SS-Net [46], BCP [32], GA-BCP [47], and ABD [33]. As shown in Table 1, incorporating our methods into MT, BCP, and ABD (i.e., Ours(MT), Ours(BCP), and Ours(ABD), they show significant improvements, which highlights the flexibility and scalability of our methods. Specifically, at a labeling data of 5%, applying our method to the basic MT model (i.e., Ours(MT)) results in performance surpassing several comparison methods. When integrated with BCP and ABD, both baseline models show stable improvements at 5% and 10% labeled data. Notably, Ours(ABD) achieves the best performance with a Dice score of 89.73% at 5% labeled data. At a 10% labeling rate, Ours(BCP) reaches a Dice score of 90.40%, achieving the best result. These results demonstrate that our bidirectional uncertainty-aware region learning strategy effectively enhances the model’s focus on difficult region learning while reducing the influence of erroneous pseudo-labels. This dual strategy improves segmentation quality and pushes the upper bound of model performance in semi-supervised settings.
Fig. 3 shows a qualitative comparison of segmentation results on the ACDC dataset, as can be observed from the figure, the baseline methods (MT, BCP, and ABD) are prone to segmentation errors in regions with complex structures or fuzzy boundaries, which are manifested as unclear boundaries or incomplete region prediction. In contrast, our method performs more robustly in these difficult regions and is able to effectively suppress erroneous predictions, presenting clearer structural boundaries and more complete region coverage.
@@ 4.4.1 ACDC dataset
Figure 3: Visualization of segmentation results on ACDC dataset with 10% labeled data. (a) Ground-truth. (b) Mean Teacher results. (c) Ours (Mean Teacher) results. (d) BCP results. (e) Ours (BCP) results. (f) ABD results. (g) Ours (ABD) results.
@@
Figure 4: Visualization of segmentation results on LA dataset with 10% labeled data. (a) Ground-truth. (b) Mean Teacher results. (c) Ours (Mean Teacher) results. (d) BCP results. (e) Ours (BCP) results.
4.4.2 LA dataset
Following previous methods, we conducted experiments on the LA dataset using 5% and 10% labeled data. We compared our method with several state-of-the-art semi-supervised segmentation approaches, including MT, SASSNet [29], DTC [30], URPC [41], MC-Net [45], SS-Net [46], BCP [32], and AllSpark [48]. As shown in Table 2, integrating our method into MT and BCP, i.e., Ours(MT) and Ours(BCP), they both achieves significant improvements, with Ours(BCP) showing the most advanced performance, demonstrating that by applying different uncertainty handling strategies to labeled and unlabeled data, it can also effectively help existing models improve their performance on 3D datasets.
We further performed visualization analysis on the LA dataset to verify the advantages of the proposed strategy in the actual segmentation results. As shown in Fig. 4, we demonstrate the segmentation results of each method from 3D view and 2D slice level, respectively. It can be observed that the baseline model is prone to prediction errors when dealing with the edges of anatomical structures or regions with complex shapes, while the segmentation performance of the model in these regions is significantly improved after integrating our proposed strategy into the baseline model.
[…]
U-Net 20% 0% 60.88 13.87
U-Net 100% 0% 91.47 84.36
CCT 10% 90% 46.73 2.14
URPC 61.92 2.94
$S$-Net 57.41 6.33
SCP-Net 55.14 19.52
BCP 79.68 7.84
ABD 81.81 1.46
Ours(BCP) 85.20 1.00
CCT 20% 80% 71.43 16.61
URPC 63.23 4.33
$S$-Net 62.31 4.36
SCP-Net 77.06 3.52
BCP 79.81 2.9
ABD 82.06 1.33
Ours(BCP) 86.01 1.04
Table 3: Comparisons with state-of-the-art semi-supervised segmentation methods on PROMISE12 dataset.
4.4.3 PROMISE12 dataset.
We also perform the comparison experiments in the PROMISE12 dataset using 10% and 20% labeled data. As shown in the Table 3, We compare our method with CCT [20], URPC [41], SS-Net [46], SCP-Net [49], BCP [32] and ABD [33]. When our strategy is integrated into the BCP framework, the proposed uncertainty-aware region learning method brings consistent and significant performance improvements. Specifically, under the 10% labeled setting, Ours(BCP) achieves the best Dice score of 85.20% and the lowest ASD of 1.00%, outperforming all other competing methods. Under the 20% labeled setting, Ours(BCP) again leads with a Dice score of 86.01% and an ASD of 1.04%. These results further confirm that our method can effectively enhance the performance of strong baseline models and push their upper performance bounds.
@@ 4.5 Ablation Studies
We conduct an ablation study to show the impact of each component in our model, including the effectiveness of the bidirectional region learning strategy, the regional learning approach, the uncertainty threshold $\mu$, the Weight in the Loss Function $\alpha$, and the impact of removing unreliable prediction regions on model learning.
[…]
@@
Table 4: Effectiveness of URL and CRL modules. “Base” means the baseline is BCP.
4.5.1 Effectiveness of Each Module in Bidirectional Region Learning
As shown in Table 4, to validate the bidirectional region learning strategies, we explore the following approaches under Ours(BCP) 10% labeled data. We demonstrate the effectiveness of each module in dual-phase region learning by progressively adding URL and CRL. When URL is added, the Dice increased to 89.45%, indicating that focusing on high-uncertainty regions in labeled data helps the model better learn from complex areas and transfer the learned knowledge to the training of unlabeled data. When CRL is added, the Dice rose to 90.20%, indicating that focusing on low-uncertainty regions in unlabeled data while reducing the learning weights of high-uncertainty regions helps mitigate the impact of erroneous pseudo-labels. By incorporating both URL and CRL into the baseline, the model has the best result, reaching 90.40% Dice and surpassing the baseline in 1.56%. These results indicate that the two modules are complementary during the training process. In this way, the model not only gains a deeper understanding of complex regions but also maintains stability and accuracy in the training of unlabeled data, thereby improving the overall performance of the model.
[…]
@@ 4.5.1 Effectiveness of Each Module in Bidirectional Region Learning
Table 5: Impact of different regional learning strategies on model performance.
@@ 4.5.2 Regional Learning Strategies
To evaluate the effectiveness of the URL and CRL modules on labeled and unlabeled data, under 10% labeled data based on Ours (BCP), we conduct different combination experiments with both modules on labeled and unlabeled data, as shown in Table 5. Through these combination strategies, we are able to analyze the effects of applying the two modules on different types of data. The experimental results indicate that all combination strategies brought significant performance improvements, demonstrating the effectiveness of the URL and CRL modules. The optimal result is achieved by applying URL on labeled data and CRL on unlabeled data. This result confirms our hypothesis: for labeled data, we should focus on learning uncertain regions. By concentrating on these uncertain areas and leveraging accurate label information for supervision, the model can gain a deeper understanding of the features in challenging regions, thus improving its segmentation ability in difficult areas. In contrast, for unlabeled data, due to the absence of ground-truth, errors in the pseudo-labels generated by the model are inevitable. Therefore, focusing on learning the deterministic regions predicted by the model can effectively reduce the interference of pseudo-label noise, thus enhancing the stability of the training. The combination of these two strategies fully exploits the different characteristics of labeled and unlabeled data, improving the model’s segmentation performance.
@@
[…]
@@
Table 6: Impact of uncertainty threshold $\mu$ settings in URL and CRL on model performance.
4.5.3 Effect of Uncertainty Thresholds $\mu$ on Model Performance
In this section, we explore the effect of the uncertainty threshold $\mu$ on model performance. Specifically, choosing the 95th percentile to identify uncertainty regions in labeled data aims to comprehensively cover areas where errors may occur and rely on accurate labels to correct them. Conversely, selecting the 99th percentile in unlabeled data aims to more accurately pinpoint uncertainty regions. As shown in Fig. 5, when the $\mu$ is set at the 95th and 97th percentiles, some uncertainty regions partially appear in the background. Generally speaking, the model can receive many background samples during training, resulting in a lower probability of prediction errors in background regions. Therefore, the choice of the 99th percentile aims to avoid misclassifying correct regions as uncertainty regions while maximizing the precision of uncertainty region localization. To validate this hypothesis, we conduct the relevant experiments, as shown in Table 6: Compared to the baseline, the combinations of $\mu$ set at 95%+95% and 99%+99% for labeled and unlabeled data both exhibit reliable performance, indirectly validating the effectiveness of the proposed module. However, their performance is still lacking compared to the 95%+99% combination. This suggests that the mixed approach of using a broader uncertainty range (95%) in labeled data and a more precise uncertainty selection (99%) in unlabeled data achieves the optimal balance. This combination effectively provides ample correction opportunities for labeled data while avoiding misjudgment of uncertain regions in unlabeled data.
Additionally, in the initial stage of training with unlabeled data, we visualize the pseudo-labels generated by the teacher model. As shown in Fig. 6, in BCP, the pseudo-labels generated by the teacher model contain many errors. However, in Ours(BCP), the pseudo-label errors generated by the teacher model trained with the 99th percentile threshold are slightly reduced. In contrast, those generated with the 95th percentile threshold are significantly fewer. This not only verifies the feasibility of our experimental setup, but also further demonstrates that focusing on learning uncontrollable regions in labeled data can enhance the model’s ability to understand and adapt to the uncontrollable regions, thus generating higher-quality pseudo-labels during the training process on unlabeled data, and improving the model’s self-supervised learning ability and overall performance.
[…]
0.2 90.05 82.36 2.25 0.62
0.5 90.40 83.01 1.63 0.41
0.7 90.26 82.76 2.28 0.58
Table 7: Impact of weighting factor $\alpha$ on model learning.
4.5.4 Weight in Loss Function
Table 7 investigates the impact of the weighting factor $\alpha$ on the loss function, evaluated on the ACDC dataset with 10% labeled data. The results show that setting $\alpha$ = 0.5 achieves the best balance, the configuration used in our experiments. Furthermore, different $\alpha$ settings outperform the original BCP baseline, indicating that our uncertainty region learning strategy remains effective even when the balance between uncertainty and certainty region learning is not optimal.
[…]